# A Unified Approach for Multi-Granularity Search over Spatial Datasets

### v3

@@ 8. Conclusions
We presented Spadas, a multi-granularity spatial data search system that seamlessly supports coarse-grained dataset search based on multiple distance metrics and fine-grained data point search. We first designed a unified two-level index, which can effectively organize spatial datasets to support multiple types of search operations, and mitigate the impact of outliers on the accuracy of similarity measures between datasets. Then, we proposed a range of pruning mechanisms, including fast bound estimation, approximation technique with error bound, and pruning in batch techniques, which can efficiently prune irrelevant datasets and data points, thereby significantly reducing the search overhead. Finally, we conducted extensive experiments using six real-world dataset collections and compared the performance of the proposed algorithms with that of related state-of-the-art algorithms. The results corroborated that the proposed system is effective, efficient, and scalable. In future work, we plan to further evaluate the effectiveness of additional distance measures.

### v5

@@ 8. Conclusions
We presented Spadas, a multi-granularity spatial data search system that seamlessly supports coarse-grained dataset search based on multiple distance metrics and fine-grained data point search. We first designed a unified two-level index, which can effectively organize spatial datasets and data points to support multiple types of search operations, and mitigate the impact of outliers on the accuracy of similarity measures. Then, we proposed a range of pruning mechanisms, including fast bound estimation, approximation technique with error bound, and pruning in batch techniques. Finally, we conducted extensive experiments using six real-world data repositories. The results corroborated that the proposed system is effective, efficient, and scalable. In future work, we plan to further evaluate the effectiveness of additional distance measures.